OpenStudy (anonymous):

Find f. (Use C for the constant of the first antiderivative and D for the constant of the second antiderivative.) f ''(x) = 4 + x^3 + x^6 i got Cx+D+2x^2+(x^5)/5+(x^8)/8 is this correct?

OpenStudy (astrophysics):

\[f'(x) = \int\limits (4+x^3+x^6) dx = 4x+\frac{ x^4 }{ 4 }+\frac{ x^7 }{ 7 } + C\] \[f(x) = \int\limits (4x+\frac{ x^4 }{ 4 }+\frac{ x^7 }{ 7 }+C) dx = \frac{ 4x^2 }{ 2 }+\frac{ x^5 }{ 20 }+\frac{ x^8 }{ 56 }+Cx+D\]

OpenStudy (astrophysics):

See your mistake?

OpenStudy (anonymous):

oh I see it thanks!!

OpenStudy (astrophysics):

Np :)